1998 – Vacuum system and EXPO 2000

Spring 1998: Progress Report

Mid of January the heating of the eastern tube started, but the phase cutting current control only survived the first half hour. The thyristors were underrated. Preparation is being made for leak testing the second arm.

The blown-up thyristors have been replaced by appropriate ones and the phase cutter works again. The eastern tube has been air baked at 200°C for about one and a half day. The vacuum bake lasted for three days. The mass 2 (hydrogen) readings of the quad in the central station showed too high values. In addition, as the quad was not baked simultaneously with the rest of the tube a lot of dirt accumulated in the quad. Another air bake and a subsequent vacuum bake (in combination with changing the gauges for clean ones) solved this problem. After switching off the heating of the tube, two leaks developed at the 600 mm COF flanges between the tube and the 600 mm gate valves. Tightening the bolts solved this problem.

The pressure in the eastern tube is now (27.3.98) 2•10-8 mbar and the quad reading shows a clear air signature. We hope it is one of the flanges in the trench and not the tube itself. As this pressure is clearly higher than the nitrogen partial pressure was before the bake, it is a new leak we are facing.

The northern tube is pumped down and has reached a pressure in the 10-5 mbar range. Leak testing is under way. The stacks for the Mode Cleaners are well on their way.

loading content
Go to Editor View